Express the number appearing in the following statements in standard form.
 Thickness of a thick paper is 0.07 mm

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Thickness of a thick paper = 0.07 mm
0.07 mm = 7/100 mm = 7/10^2= 7×10^-2mm

Thanks